OpenStudy (anonymous):

Is my answer correct? ;) What value is added to both sides of the equation x2 - 8x = -10 in order to solve by completing the square? Answer: I know it would be 16 if the answer to the problem was -11, but in this case, the answer I got was "-8"

jimthompson5910 (jim_thompson5910):

what do you mean by "I know it would be 16 if the answer to the problem was -11"

OpenStudy (anonymous):

I applied a certain step, and I figured I'd do the opposite for this problem....

jimthompson5910 (jim_thompson5910):

and what step was that

OpenStudy (anonymous):

a^2-2ab+b^2 = (a-b)^2 a = x => 2ab = 8x => b = 8x/2x => b = 4 => b^2 = 16 16 is to be added to both sides of the equation.

jimthompson5910 (jim_thompson5910):

good, you can also use this shortcut x coefficient: -8 cut the x coefficient in half: -8/2 = 4 square that last result: 4^2 = 16

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

that makes it easier, thanks :)